# Analyzing Games in Maker Protocol Part One: A Multi-Agent Influence Diagram Approach Towards Coordination

### v1

@@
Abstract
Decentralized Finance (DeFi) ecosystems, exemplified by the Maker Protocol, rely on intricate games to maintain stability and security. Understanding the dynamics of these games is crucial for ensuring the robustness of the system. This motivating research proposes a novel methodology leveraging Multi-Agent Influence Diagrams (MAID), originally proposed by Koller and Milch, to dissect and analyze the games within the Maker stablecoin protocol. By representing users and governance of the Maker protocol as agents and their interactions as edges in a graph, we capture the complex network of influences governing agent behaviors. Furthermore in the upcoming papers, we will show a Nash Equilibrium model to elucidate strategies that promote coordination and enhance economic security within the ecosystem. Through this approach, we aim to motivate the use of this method to introduce a new method of formal verification of game theoretic security in DeFi platforms.
Keywords:
DeFi Multi-Agent Influence Diagrams Maker ProtocolGame Theory Formal Verification Security
1 Introduction
Decentralized Finance (DeFi) [2] has emerged as a transformative force in the financial landscape, offering innovative solutions that challenge traditional centralized systems [1]. At the forefront of this movement lies the Maker Protocol [4], a cornerstone of the DeFi ecosystem, providing the infrastructure for the creation and governance of the Dai stablecoin (see [5]). However, the stability and security of DeFi platforms like Maker heavily rely on intricate coordination games, which necessitate a deep understanding of their dynamics.
In this research, we propose a novel methodology for dissecting and analyzing the games within the Maker Protocol and in turn, similar stablecoin protocols, leveraging Multi-Agent Influence Diagrams (MAIDs) [3]. Originally introduced by Koller and Milch in 2003, MAIDs offer a powerful framework for modeling complex decision-making scenarios involving multiple agents with competing or aligned objectives. By representing users of the Maker protocol as agents and their interactions as edges in a graph, we capture the intricate network of influences that govern user behaviors within the ecosystem.
Our research aims to shed light on the interplay between incentives, behaviors, and outcomes within the Maker Protocol. By employing MAIDs, we aim to systematically analyze the incentives driving user decisions, identify vulnerabilities in the system, and optimize incentive mechanisms to enhance the stability and security of the platform. Furthermore, we will introduce a Nash Equilibrium model [6, 7] in next papers to elucidate strategies that promote coordination among users, fostering economic security and resilience in the DeFi ecosystem.
Through this approach, we not only contribute to a deeper understanding of the Maker Protocol but also provide a methodology that can be extended to other DeFi platforms. By fortifying coordination security and mitigating risks, our research seeks to pave the way for the continued growth and adoption of decentralized financial systems, ushering in a new era of financial inclusivity and innovation. This paper focuses on construction of MAIDs for the Maker protocol and the subsequent iterations would involve employing the ideas of MAID to conduct a full length equilibrium analysis of the protocol using the mathematical ideas presented in the following sections.
2 Background
2.1 Multi Agent Influence Diagrams
@@
Multi Agent Influence Diagrams (MAIDs) combine the ideas of Bayesian Networks (BNs) [8] and Influence Diagrams [9] to describe decision problems related to multiple agents. MAIDs are graphical models used to represent decision-making scenarios involving multiple agents, each with their own objectives, beliefs, and actions. They extend the traditional influence diagram framework to accommodate the complexities of interactions and dependencies among multiple decision-makers. In a MAID, nodes represent variables such as decisions, uncertainties, utilities, and influences, while directed edges indicate causal relationships or dependencies between them. Each agent is associated with a subset of nodes, representing their local decisions, beliefs, and preferences (see [3]). MAIDs allow for the explicit representation of interactions between agents, including cooperation, competition, coordination, and negotiation. This enables the analysis of strategic interactions and the prediction of outcomes resulting from the decisions of multiple agents. There has been a lot of related work done by Google Deepmind (see [10, 11]) on agent incentives for artificial intelligence architectures which forms the underlying motivator to try and bring the same ideas to DeFi.,
2.2 Maker Protocol
The Maker Protocol operates as a decentralized finance (DeFi) platform on the Ethereum blockchain, facilitating the creation of Dai, a stablecoin pegged to the US dollar. As per [4], Users engage with the protocol by depositing cryptocurrency assets, such as Ethereum or other tokens, into smart contracts as collateral. Based on the value of this collateral, users can then generate Dai, typically up to a predetermined percentage of the collateral’s worth. To maintain the stability and security of the system, the protocol enforces various risk parameters, including collateralization ratios and debt ceilings, which dictate the relationship between collateral and the amount of Dai that can be generated. Users pay stability fees on the Dai they generate, serving as interest rates that contribute to the protocol’s stability. Governance of the Maker Protocol is decentralized, with MKR token holders participating in decision-making through a voting mechanism. These holders vote on proposals to adjust risk parameters and other protocol settings, ensuring the protocol remains responsive to market conditions and emerging risks. Additionally, the protocol includes mechanisms for liquidating collateral if its value falls below a certain threshold, helping to mitigate systemic risk and maintain the stability of the Dai ecosystem. Overall, the Maker Protocol offers users a decentralized and trustless means of generating a stable cryptocurrency, providing access to the benefits of DeFi while minimizing counterparty risk.
3 Model
3.1 Mathematical Exposition
As per Koller and Milch [3], to define a Multi-Agent Influence Diagram (MAID), we start with a collection of agents. We represent them as $\textbf{A}=\{A_{i}\}$ for $i\in N$. These agents interact within the Maker Protocol, which is represented by a set of chance variables denoted as $\chi$, and a set of decision variables $D_{A_{i}}$ specific to each agent $A_{i}\in A$.
Chance variables represent uncertain events or outcomes, similar to nodes in Bayesian networks[8], while decision variables are the choices that each agent $A_{i}$ can make, depicted as rectangular nodes in the diagram. We use $D$ to denote $\cup_{A_{i}\in A}D_{A_{i}}$ The agents’ preferences and objectives are captured using utility variables, denoted as $U_{A_{i}}$ for each agent $A_{i}\in A$, where each agent has a set of utility variables. Similar to the decision variables, we use $U$ to denote $\cup_{A_{i}\in A}U_{A_{i}}$. These utility variables quantify the agent’s satisfaction or payoff based on the outcomes of the decision and chance variables. In the Maker Protocol context, chance variables represent factors like market conditions and collateral fluctuations, while decision variables denote actions taken by users and governance entities.
Mathematically, a MAID is represented as a directed acyclic graph (which is used to denote causal relationships, see [12]) where each variable corresponds to a node, and the relationships between variables are depicted through directed edges. For each chance variable $X_{i}\in\chi$, the MAID specifies a conditional probability distribution (CPD) given the instantiations of its parent variables $Pa(N)\subset\chi\cup D$ for each node $N$ in the diagram, which is $P(N|Pa(N))$. Similarly, for each utility variable $U_{A_{i}}$, the MAID specifies a conditional probability distribution, ensuring that the utility value is deterministic based on the values of its parents. The total utility derived by an agent from a particular instantiation of the environment is computed as the sum of the utility values of all relevant utility variables.
Agents determine their actions at decision nodes based on the variables they observe prior to making a decision, typically represented by the parent variables of the decision node. The decision-making process is defined by a set of decision rules, which specify the agent’s strategy or behavior at each decision node. In the context of the Maker Protocol, users and governance entities make decisions based on observed market conditions, collateral fluctuations, and similarly relevant factors to optimize their utility. By utilizing MAIDs, stakeholders can model and analyze the complex interactions and decision-making processes within the Maker Protocol ecosystem to enhance stability, security, and efficiency as we will demonstrate in the next section. The next portion describes the definitions and theorems, from Koller and Milch’s paper [3], which will be useful in our analysis. Readers are encouraged to refer to the original paper to know more about the ideas in detail.
Definition 1
@@ 3.1 Mathematical Exposition
A decision rule for a decision variable $D$ is a function that maps each instantiation $pa$ of parents $\mathcal{P}(D)$ to a probability distribution over the domain $\text{dom}(D)$. An assignment of decision rules to every decision $D_{A_{i}}$ for a particular agent in $A$ is called a strategy.
We define the payoffs at the leaves by computing a distribution over the utility nodes, given an instantiation of $\theta$. For a leaf $\ell$, the payoff for agent $A$ is:
$u_{A}(\ell)=\mathbb{E}_{\theta}[u_{A}\,|\,\text{inst}_{\ell}]$
Theorem 3.1

### v2

# Multi Agent Influence Diagrams for DeFi Governance

@@
Abstract.
Decentralized Finance (DeFi) governance models have become increasingly complex due to the involvement of numerous independent agents, each with their own incentives and strategies. To effectively analyze these systems, we propose using Multi Agent Influence Diagrams (MAIDs) (Koller and Milch, 2001) as a powerful tool for modeling and studying the strategic interactions within DeFi governance. MAIDs allow for a comprehensive representation of the decision-making processes of various agents, capturing the influence of their actions on one another and on the overall governance outcomes. In this paper, we study a simple governance game that approximates real governance protocols and compute the Nash equilibria using MAIDs. We further outline the structure of a MAID in MakerDAO.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Decentralized Finance (DeFi) governance protocols empower token holders to actively participate in shaping the future of the platform by voting on a variety of important decisions. These decisions can range from technical adjustments, such as parameter updates, to broader protocol improvements that impact the entire ecosystem. The process of deciding how to vote on a particular proposal is complex and influenced by multiple factors. Firstly, voter behavior is often shaped by the actions and decisions of other voters. This phenomenon, known as social influence or herding behavior, can significantly impact the outcome of votes as individuals may be swayed by the perceived consensus or the actions of influential stakeholders within the network. Secondly, external factors, such as market sentiment, play a critical role in shaping voter decisions. Market trends, news events, and the overall economic climate can affect how token holders perceive the implications of a proposal and thus influence their voting behavior. Understanding the dynamics of these voting games is crucial for ensuring the robustness of the system. For the analysis of such games, we propose to use Multi Agent Influence Diagrams (MAIDs). MAIDs combine the ideas of Bayesian Networks (BNs) (Pearl, 2014) and Influence Diagrams (Howard and Matheson, 2005) to describe decision problems related to multiple agents. MAIDs are graphical models used to represent decision-making scenarios involving multiple agents, each with their own objectives, beliefs, and actions. They extend the traditional influence diagram framework to accommodate the complexities of interactions and dependencies among multiple decision-makers. In a MAID, nodes represent variables such as decisions, uncertainties, utilities, and influences, while directed edges indicate causal relationships or dependencies between them. Each agent is associated with a subset of nodes, representing their local decisions, beliefs, and preferences (see (Koller and Milch, 2001)). MAIDs allow for the explicit representation of interactions between agents, including cooperation, competition, coordination, and negotiation. This enables the analysis of strategic interactions and the prediction of outcomes resulting from the decisions of multiple agents. There has been a lot of related work done by Google Deepmind (see (Everitt et al., 2021, 2019)) on agent incentives for artificial intelligence architectures which forms the underlying motivator to try and bring the same ideas to DeFi. Just as Bayesian networks make explicit the dependencies between probabilistic variables, MAIDs make explicit the dependencies between the probabilistic variables that determine the decision taken by each agent in the game. Note that for every MAID, there is a corresponding extensive form game.
The paper is structured as follows. The next section introduces MAIDs. Section 3 describes a simple governance model. In Section 4, we solve the MAID for Nash equilibria in a single-agent and in a multi-agent setting. Section 5 presents a MAID for MakerDAO governance. Finally, Section 6 concludes.
2. Multi Agent Influence Diagrams (MAIDs)
In this section, we recall the necessary notions from Koller and Milch (Koller and Milch, 2001) to introduce MAIDs and the algorithm to compute Nash equilibria of MAIDs.
Definition 2.1 (Chance, Decision, and Utility Variables).
We define the variables as follows:
  • A chance variable represents uncertain events or outcomes, similar to nodes in Bayesian networks (Pearl, 2014) that affect each agent’s decisions. The set of all chance variables is denoted by $\chi$.
  • A decision variable is a variable whose value is chosen by an agent and these are the decisions that each agent $a\in\mathcal{A}$ can make. The set of all chance variables is denoted by $\mathcal{D}$.
  • A utility variable specifies a utility function for each agent $a\in\mathcal{A}$. The set of all utility variables is denoted by $\mathcal{U}$.
We define $\mathcal{V}:=\chi\cup\mathcal{D}\cup\mathcal{U}$ the set of all variables.
Definition 2.2 (MAID).
A MAID is tuple $\mathcal{M}=\langle\mathcal{A},\chi,\mathcal{D},\mathcal{U}\rangle$, where $\mathcal{A}$ is the (finite) set of agents, $\chi$ the (finite) set of chance variables, $\mathcal{D}$ the (finite) set of decision variables, and $\mathcal{U}$ the (finite) set of utility variables. For each agent $a\in\mathcal{A}$ the individual sets of variables are $\langle\mathcal{D}_{a},\mathcal{U}_{a}\rangle$ and thus $\mathcal{D}=\bigcup_{a\in\mathcal{A}}\mathcal{D}_{a},\mathcal{U}=\bigcup_{a\in\mathcal{A}}\mathcal{U}_{a}$. Note that $\mathcal{M}$ defines a directed acyclic graph (DAG) (Williams et al., 2018) structure where variables are nodes.
Definition 2.3 (Causal Influence Diagram).
A Causal influence diagram (CID) (Everitt et al., 2021) for an agent $a\in\mathcal{A}$ is a tuple $(\mathcal{V}_{a},E_{a})$ where $(\mathcal{V}_{a},E_{a})$ is a directed acyclic graph (DAG) with a set of vertices $\mathcal{V}_{a}=\chi\cup\mathcal{D}_{a}\cup\mathcal{U}_{a}$ connected by directed edges $E_{a}\subseteq\mathcal{V}_{a}\times\mathcal{V}_{a}$.
Definition 2.4 (Parent Set).
Let $\mathcal{M}=\langle\mathcal{A},\chi,\mathcal{D},\mathcal{U}\rangle$ be a MAID. For every variable $X\in\mathcal{V}$, $\mathtt{Pa}(X)\subset\mathcal{V}\setminus\mathcal{U}$ is the set of all parent nodes. By definition (and in keeping with the structure of a DAG) utility variables do not have child nodes. For any agent $a\in\mathcal{A}$, the parent set $\mathtt{Pa}(D)$ for $D\in\mathcal{D}_{a}$, is the set of variables whose values agent $a$ knows when he chooses a value for D.
Definition 2.5 (Conditional Probability Distribution (CPD)).
Let $\mathcal{M}$ be a MAID. Then, for each chance variable $X\in\chi$, $\mathcal{M}$ induces a conditional probability distribution (CPD) $\mathbb{P}(X|\mathbf{pa})$ for each instantiation (i.e for each instance of variable) $\mathbf{pa}\in\mathtt{Pa}(X)$. Similarly, for each utility variable $U\in\mathcal{U}$, $\mathcal{M}$ induces a CPD $\mathbb{P}(U|\mathbf{pa})$ for each instantiation $\mathbf{pa}\in\mathtt{Pa}(X)$.111there are more things on the utility. maybe we need to add them.
Definition 2.6 (Decision Rule, Strategy).
@@ 2. Multi Agent Influence Diagrams (MAIDs)
A decision rule for a decision variable $D\in\mathcal{D}$ is a function that maps each instantiation $\mathbf{pa}$ of parents $\mathtt{Pa}(D)$ to a probability distribution over the domain $\mathtt{dom}(D)$. An assignment of decision rules to every decision $D\in\mathcal{D}_{a}$ for a particular agent in $a\in\mathcal{A}$ is called a strategy.
Definition 2.7 ((Partial) Strategy Profile).
An assignment $\sigma$ of decision rules to every decision $D\in\mathcal{D}$ is called a strategy profile. A partial strategy profile $\sigma_{\mathcal{E}}$ is an assignment of decision rules to a subset $\mathcal{E}\in\mathcal{D}$.
Remark 1.
Given a MAID $\mathcal{M}$, then a partial strategy profile $\sigma_{\mathcal{E}}$ induces a new MAID $\mathcal{M}{[\sigma_{\mathcal{E}}]}$ where the elements of $\mathcal{E}\subset\mathcal{D}$ are the chance variables and for each $D\in\mathcal{E}$, $\sigma_{\mathcal{E}}(D)$ is a CPD.
Definition 2.8 (Joint Distribution).
If $\mathcal{M}$ is a MAID and $\sigma$ is a strategy profile for $\mathcal{M}$, then the joint distribution for $\mathcal{M}$ induced by $\sigma$, denoted $P_{\mathcal{M}[\sigma]}$, is the joint distribution over $\mathcal{V}$ defined by the Bayes network where:
  • the set of variables is $\mathcal{V}$;
  • for $X,Y\in\mathcal{V}$, there is an edge $X\rightarrow Y$ iff $X\in\mathtt{Pa}(Y)$;
  • for all $X\in\mathcal{V}\setminus\mathcal{D}$, the CPD for $X$ is $\mathbb{P}(X)$;
  • for all $D\in\mathcal{D}$, the CPD for $D$ is $\sigma(D)$.
We are ready to motivate the definition of agent utility.
Definition 2.9 (Expected Utility).
The expected utility that any agent $a\in\mathcal{A}$ anticipates in a MAID $\mathcal{M}$ when the agents play strategy profile $\sigma$ is
(1) $EU_{a}(\sigma)$ $=\sum_{(u_{1},u_{2},\ldots,u_{m})\in dom(\mathcal{U}_{a})}P_{\mathcal{M}[\sigma]}(u_{1},u_{2},\ldots,u_{m})\sum_{i=1}^{m}u_{i}$
(2) $=\sum_{U\in\mathcal{U}_{a}}\sum_{u\in dom(U)}P_{\mathcal{M}[\sigma]}(U=u)\cdot u$
Definition 2.10 (Optimal Strategy).
Let $\mathcal{E}$ be a subset of decision variables $\mathcal{D}_{a}$, and let $\sigma$ be a strategy profile. We say that $\sigma^{*}_{\mathcal{E}}$ is optimal for the strategy profile $\sigma$ if, in the induced MAID $\mathcal{M}[\sigma_{\mathcal{E}}]$, where the only remaining decisions are those in $\mathcal{E}$, the strategy $\sigma_{\mathcal{E}}$ is optimal.
Formally, this means that for all strategies $\sigma^{\prime}_{\mathcal{E}}$:
$EU_{a}(\sigma_{-\mathcal{E}},\sigma^{*}_{\mathcal{E}})\geq EU_{a}(\sigma_{-\mathcal{E}},\sigma^{\prime}_{\mathcal{E}}).$
Definition 2.11 (Nash Equilibrium for a MAID).
A strategy profile $\sigma$ is a Nash equilibrium for a MAID $\mathcal{M}$ if for all agents $a\in\mathcal{A}$, $\sigma_{\mathcal{D}_{a}}$ is optimal for the strategy profile $\sigma$.
Now that we understand the Nash equilibrium in a MAID, we present an algorithm that helps us compute the Nash Equilibria (Nash et al., 1950).
Proposition 2.12 (MAID and Extensive-Form Game (Koller and Milch, 2001)).
Let $\mathcal{M}$ be a MAID. Then, there is a corresponding extensive-form game tree $\mathcal{T}$. For any strategy profile $\sigma$, the payoff vector for $\sigma$ in $\mathcal{M}$ is the same as for $\sigma$ in $\mathcal{T}$.
2.1. Algorithm for Computing the Nash Equilibrium
In order to compute the Nash equilibrium for a particular MAID $\mathcal{M}$, we will require the use of a few more definitions as follows:
Definition 2.13 (s-reachable).
A node $D^{\prime}$ is $s$-reachable from a node $D$ in a MAID $\mathcal{M}$ if there exists some utility node $U\in\mathcal{U}_{D}$ such that adding a new parent $D^{\prime\prime}$ to $D^{\prime}$ would create an active path in $\mathcal{M}$ from $D^{\prime\prime}$ to $U$, given the evidence set $\mathtt{Pa}(D)\cup\{D\}$.
Here, a path is considered active in a MAID if it is active in the same graph viewed as a Bayesian network (BN) (Pearl, 2014).
Definition 2.14 (Relevance Graph).
The relevance graph for a MAID $\mathcal{M}$ is a directed graph whose nodes are the decision nodes of $\mathcal{M}$. There is an edge directed from node $D$ to node $D^{\prime}$ (denoted $D\rightarrow D^{\prime}$) if and only if $D^{\prime}$ is $s$-reachable from $D$ in $\mathcal{M}$.
Definition 2.15 (Strongly Connected Component).
A set $\mathcal{S}$ of nodes in a directed graph is a strongly connected component (SCC) if for every pair of nodes $D\neq D^{\prime}\in\mathcal{S}$, there exists a directed path from $D$ to $D^{\prime}$. A maximal SCC is an SCC that is not a strict subset of any other SCC.
We can find the maximal SCCs of a relevance graph in linear time, by constructing a component graph (Cormen et al., 2022) whose nodes are the maximal SCCs of the graph. There is an edge from component $\mathcal{C}_{i}$ to component $\mathcal{C}_{j}$ in the component graph if and only if there is an edge in the relevance graph from some element of $\mathcal{C}_{i}$ to some element of $\mathcal{C}_{j}$.
The component graph is always acyclic, so we can define an ordering $C_{1},C_{2},\cdots C_{m}$ over the SCCs, such that whenever $i<j$ no element of $C_{i}$ is $s$-reachable from any element of $C_{j}$.
Definition 2.16 (Topological Ordering (Cormen et al., 2022)).
Let $(\mathcal{V},E)$ be a DAG. A topological ordering of $(\mathcal{V},E)$ is a linear ordering of its vertices such that for every directed edge $(u,v)$ in $E$, vertex $u$ comes before $v$ in the ordering.
Next, we present the algorithm for computing the Nash equilibria for a MAID.
Algorithm 1 Computing Nash Equilibria (Koller and Milch, 2001)
Given a MAID $\mathcal{M}$
a topological ordering $C_{1},C_{2},\cdots C_{m}$ of the component graph derived from the relevance graph for $\mathcal{M}$
Let $\sigma^{0}$ be an arbitrary fully mixed strategy profile
for $i=0$ through $m-1$do
     Let $\tau$ be a partial strategy profile for $C_{m-i}$ that is a Nash equilibrium in $\mathcal{M}[{\sigma^{i}}_{-C_{m-i}}]$
     Let $\sigma^{i+1}=({\sigma^{i}}_{-C_{m-i}},\tau)$
end for
Output $\sigma^{m}$ as an equilibrium of $\mathcal{M}$
The algorithm processes each Strongly Connected Component (SCC) in reverse order, determining an equilibrium strategy profile for each SCC in the Multi-Agent Influence Diagram (MAID). This is done based on the previously selected decision rules, with arbitrary decision rules assigned to decisions that are not pertinent to the current SCC. In this derived MAID, the only decision nodes that remain are those within the current SCC, while all other decision nodes are treated as chance nodes. To find the equilibrium in this transformed graph, a subroutine designed to identify equilibria in games is used. The game is then transformed into a game tree, and a standard game-solving algorithm is applied to identify a decision rule that maximizes the expected utility for the individual agent.
Before stating the result of correctness of the algorithm, we need the following definition.
Definition 2.17 (Perfect Recall).
An agent $a\in\mathcal{A}$ has perfect recall w.r.t. a total order $D_{1},\ldots,D_{n}$ over $\mathcal{D}_{a}$ if for all $D_{i},D_{j}\in\mathcal{D}_{a}$, $i<j$ implies that $D_{i}\in\mathtt{Pa}(D_{j})$ and $\mathtt{Pa}(D_{i})\subset\mathtt{Pa}(D_{j})$.
Proposition 2.18 (Correctness of Algorithm 1 (Koller and Milch, 2001)).
Suppose every agent has perfect recall. Then, the output $\sigma^{m}$ of Algorithm 1 is a Nash equilibrium for $\mathcal{M}$.
3. Model
Let us first describe a simple governance game. Assuming that there are $n$ homogeneous agents (meaning that all of these agents exhibit same exact behaviour given similar situations) who own stake in the governance of a DeFi protocol. Assume that the DeFi protocol has a governance token \$GOV and each of these agents owns some amount of \$GOV that would enable them to vote on proposals. Proposals can be brought forward by members of the governance for the rest to make a decision on. It is assumed that the member who brings in the proposal will choose to remain neutral and will abstain from voting. For example, in a DeFi lending borrowing protocol the proposal could be to strengthen the risk parameters of the collateral assets to allow a user to borrow less amount of an asset at the same amount of collateral simply to preserve the sustainability of the protocol. Let us visualise this model through the lens of MAIDs. Here we define the input variables of the MAID $\mathcal{M}=(\mathcal{A},\chi,\mathcal{D},\mathcal{U})$ in the following way:
  1. (1)
    $\mathcal{A}$: set of $n$ agents;
  2. (2)
    $\mathcal{D}_{a}=\{d_{a}\}$: set of decision variables for each agent $a\in\mathcal{A}$, where $d_{a}=\{\texttt{yes},\texttt{no}\}$ (decisions about a proposal);
  3. (3)
    $\mathcal{U}_{a}=\{u_{a}\}$: set of utility variables for each agent $a\in\mathcal{A}$ (How sustainable this decision would be for the protocol? Will it benefit the protocol?)
  4. (4)